# Physical aspects of the deformation of $\mathds{Z}_{2}$ kinks in a generalized $\phi^{4}$ model

### v1

@@ Abstract
We study in a generalized $\phi^{4}$ model the formation of BPS kink/antikink configurations with compacton-like profiles. One observes that the positive parameter controlling the generalizing function promotes an infinity degenerescence of the BPS solutions. Furthermore, the analysis of the excitation spectrum of the model shows that for values of the generalizing parameter closest to zero, there are both translational and vibrational modes. On the other hand, for sufficiently large values of the generalizing parameter, one does not observe the zero mode, and only vibrational modes emerge. Thus, the emergence of bound states of solitons (bions) and resonance phenomena is guaranteed when analyzing the scattering of kink/antikink structures. In this way, one notes that depending on the initial velocity, the collision can be inelastic or quasi-elastic, even in the case of compacton-like configurations. Finally, we apply the Differential Configurational Complexity technique to obtain the parameter values providing the most likely field profiles.
@@ I Introduction
A second physical aspect of our study is the scattering process of the compacton-like structures. Since the 1970s [19, 20], the phenomenon of kink scattering and similar structures has attracted enough attention from the academic community. Among the techniques developed to analyze this phenomenon, we mention the analytical collective coordinate approach [21, 22, 23, 24], and also the numerical methods have recently become a powerful tool for studying the dynamics of such phenomenon [25, 26, 27]. For instance, resonance phenomena, such as escape windows and quasi-resonances, have been discovered through numerical methods of kink dispersion. In this context, Ref. [8] shows the compacton-like profiles’ collision process but does not report the appearance of configurations called bions [^1] The bions [28] are “temporary bound-states” formed during the kink-antikink collision process, and they possess a topological charge, a combination of the topological charges of the original configurations.. Thus, we will seek structures emergent during the scattering process of the generalized kinks found by us.
We finalize our study by employing arguments derived from Configurational Entropy (CE) [29, 30] to delineate the most likely structures that may emerge in the model. Specifically, to reach our objective, one uses a variant of the CE called Differential Configurational Complexity (DCC) [31] that quantifies the complexity underlying the construction of a configuration in a field model [32, 33, 34, 15].
The manuscript is delineated as follows to present our analysis. In Section II, we study the BPS structure of a (1+1)-dimensional generalized scalar field model and the emerging structures endowed with BPS properties. The principal aim of our inquiry is to analyze the transition of the kink solutions towards a compacton-like profile. In Sec. III, one analyzes both the excitation spectrum and the scattering phenomena of the solutions found in the previous section. In Sec. IV, one presents a study on the DCC, seeking to discern the most likely field configurations inherent to the proposed theory. Finally, we present our remarks and conclusions in Sec. V.
@@ II.1 Modifying the $\phi^{4}$ model
To continue our study, we need to choose both the superpotential $W(\phi)$ and the generalizing function $f(\phi)$ to define a BPS potential (6) that preserves the $\mathds{Z}_{2}$ [^3] A system possesses a $\mathds{Z}_{2}$ symmetry when the action is invariant under the transformation $\phi\rightarrow-\phi$. symmetry [6, 36] but at the same time able to promote the spontaneous symmetry breaking of it. This way, for our purpose, we use a $\phi^{4}$ model defined by the following superpotential,
@@
As a consequence of the presence of the generalized function, as shown below, one has the emergence of a stability potential type infinity well, localizing the field fluctuations at a finite space.
III Excitation spectra and scattering of structures
@@ III.1 Excitation spectra
We display the stability potential (27) in Figs. 5, 7, and 8.
Utilizing the finite element method, we investigate the model eigenstates [19, 41]. We first expose in Figs. 5 and 6 the stability potential and corresponding eigenfunctions of the case $m=0$ (i.e., those of the standard $\phi^{4}$-theory). In this scenario, the model admits a translational mode (or zero mode, i.e., the eigenfunction with null eigenvalue, $\omega=0$) and one vibrational eigenstate ($\omega^{2}>0$). For $m\neq 0$, one notes through numerical inspection that this spectrum remains unchanged for small $m$ values (i.e., up to values of $m$ on the order of $10^{-2}$). However, the translational mode disappears for larger $m$ values, so only vibrational modes emerge. The stability potential and the respective first vibrational mode (ground state) with its eigenvalue associated are depicted in figures 7[(a) and (b)] and 8[(a) and (b)]. Notably, the presence of vibrational modes suggest the occurrence of resonances in the scattering of configurations with opposing topological charges.
@@
Refer to captionRefer to caption
@@
Figure 7: (a) Stability potential $U(x)$ (27) keeping $\lambda=1$ and $m$ varying. (b) Eigenfunctions of Eq. (22) representing the ground states (first vibrational modes) for some values of $m$ and $\lambda=1$
Refer to captionRefer to caption
(a)                                                             (b) .
Figure 8: (a) Stability potential $U(x)$ (27) keeping $m=0.45$ and $\lambda$ varying. (b) Eigenfunctions of Eq. (22) representing the ground states (first vibrational modes) for some values of $\lambda$ and $m=0.45$
Refer to captionRefer to caption
@@
The results obtained here show, for values of $m$ sufficiently greater than zero, the existence alone of vibrational mode(s), indicating that when configurations tend towards a compact-like profile, the resonance phenomenon will occur in kink-antikink scattering. Similarly, one emphasizes that experimental studies on kink collisions have also observed resonance phenomena even in systems without vibrational modes [19]. Furthermore, we are aware of two possible collision types for the system: the first produces energy emission with the presence of the resonance phenomenon, and the second is a collision in which the kink interactions form bions along the time evolution and energy dissipation [8, 43].
@@ III.2 Scattering of structures
Figures 11 and 12 show the analysis of the collisions of the kink-antikink structures for several initial velocity values. One notes that also, throughout the scattering, kink/antikink solutions behave like compacted configurations when $m\to 0.5$ and $\lambda\to\infty$. Besides, it exhibits critical velocities ($v_{\text{cr}}$) that alter the scattering profile. For instance, for $v_{\text{in}}<v_{\text{cr}}^{(1)}\simeq 0.15$, it stays evident that during the temporal evolution, the kink-antikink collide, annihilating each other and radiating energy. Conversely, one obtains an inelastic collision for an initial velocity whose value belongs to the interval <0.15,0.25]<0.15,0.25]< 0.15 , 0.25 ]. On the other hand, for $v_{\text{in}}>0.25$, the compacton-like configurations, here generated, experiment quasi-elastic collisions; see Fig. 13. Thus, the compacton-like configurations in the external scattering process support both inelastic and quasi-elastic collisions, and they manifest as solitary waves with physical characteristics akin to kinks.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
By looking closest, the numerical results shown by Figs. 11, 12, and 13, it is noted that for initial velocities smaller than $0.25$, the $K\bar{K}$ pair interacts forming oscillations (bions [^4] Bions arise during the collision of kinks and have interesting characteristics, such as topological charges, which are a combination of the topological charges of the original kinks. One can interpret the bions as a pair of “temporary kinks” that form during a collision.).For instance, when $v_{\text{in}}=0.10$, the first oscillation emerges. These oscillations increase as the velocity increases, regardless of the values of $m$ or $\lambda$. However, upon reaching the value of $v_{\text{in}}=0.25$, as time evolution, only three oscillations will occur, and then the $K\bar{K}$ pair scatters with the resonance phenomenon (oscillation amplitudes between $0.8$ and $1.2$) present for a time range greater than $200$. Finally, one notes that when $v_{\text{in}}\geq 0.35$, regardless of the value of $\lambda$, there will be a quasi-elastic scattering process with extreme regions of the $K\bar{K}$ pair undergoing oscillations (resonance) between 00 and $1.2$
IV The Differential Configurational Complexity (DCC)
[…]
Now, let us consider the numerical solutions depicted in Figs. 2 and 3, and employing a numerical approach [^6] The numerical method applied was the interpolation method. This method is used with steps of $10^{-4}$ for the range $[-1000,1000]$., we compute the integrals (31) and (33). We expose the numerical solutions obtained in Figs. 14(a) and 14(b).
Refer to captionRefer to caption
(a)                                                                 (b)
@@ V Final remarks
Analyzing the excitation spectrum of the model for fixed $\nu$ and $\lambda$ and for small values of $m$, we find a discrete spectrum with translational and vibrational modes. On the other hand, for values of $m$ sufficiently larger than zero, the excitation spectrum does not have a zero mode and possesses vibrational modes alone. Consequently, the resonance phenomena are guaranteed to occur during the scattering process of configurations with opposite topological charges. This way, the study reveals that the scattering of the compacton-like solutions presents the following features depending on the value of the initial velocity ($v_{\text{in}}$): for $v_{\text{in}}\simeq 0.15$ they collide and annihilate each other by radiating energy. On the other hand, for an initial velocity occurring in the interval $[0.15,0.25]$, an inelastic collision happens. These results reveal that the compacton-like configurations behave as solitonic waves with physical aspects similar to the ones arising in the kink/antikink collision processes. Furthermore, the compacton-like solutions collide quasi-elastically when the initial velocity is $v_{\text{in}}>0.25$.
Finally, we conclude our study by applying the DCC technique to determine the most suitable values for the $m$ parameter of the function $f(\phi)=\cos(m\pi\phi)$ and for the parameter $\lambda$. Thus, through numerical inspection, the DCC leads us to conclude that field configurations with lower complexity are more likely to emerge when $\lambda\simeq 300$ and $m\to 0.5$, i.e., compacton-like structures. It is a significant result because, despite infinite degeneracy, values of $m$ and $\lambda$ exist, describing the best configuration possible for the system.

### v2

# Physical aspects of the deformation of $\mathbb{Z}_{2}$ kinks in a generalized $\phi^{4}$ model

@@ Abstract
We study a generalized $\phi^{4}$ model that gives rise to BPS kink/antikink configurations with compacton-like profiles. One observes that the positive parameter controlling the generalizing function promotes an infinity degenerescence of the BPS solutions. We then use the Differential Configurational Complexity technique to distinguish the degenerate configurations, which allows us to obtain the parameter values providing the most likely field profiles. Besides, the analysis of the excitation spectrum of the model shows the existence of translational and vibrational modes. Thus, the emergence of bound states of solitons (bions) and resonance phenomena is guaranteed when analyzing the scattering of kink/antikink structures. In this way, one notes that depending on the initial velocity, the collision can be inelastic or quasi-elastic, even in the case of compacton-like configurations.
@@ I Introduction
A second analysis concerns employing arguments derived from Configurational Entropy (CE) [29, 30] that allow us to distinguish and delineate the most likely structures that may emerge in the model. Specifically, to reach our objective, one uses a variant of the CE called Differential Configurational Complexity (DCC) [31] that quantifies the complexity underlying the construction of a configuration in a field model [32, 33, 34, 15].
We finalized our study by discussing the scattering process of compacton-like structures. Since the 1970s [19, 20], the phenomenon of kink scattering and similar structures has attracted enough attention from the academic community. Among the techniques developed to analyze this phenomenon, we mention the analytical collective coordinate approach [21, 22, 23, 24], and also the numerical methods have recently become a powerful tool for studying the dynamics of such phenomenon [25, 26, 27]. For instance, resonance phenomena, such as escape windows and quasi-resonances, have been discovered through numerical methods of kink dispersion. In this context, Ref. [8] shows the compacton-like profiles’ collision process but does not report the appearance of configurations called bions [^1] The bions [28] are “temporary bound-states” formed during the kink-antikink collision process, and they possess a topological charge, a combination of the topological charges of the original configurations.. Thus, we will seek structures emergent during the scattering process of the generalized kinks found by us.
The manuscript is delineated as follows: Section II studies the BPS structure of a (1+1)-dimensional generalized scalar field model and the emerging structures endowed with BPS properties. The principal aim of our inquiry is to analyze the transition of the kink solutions towards a compacton-like profile. In Sec. III, one presents a study on the DCC, seeking to discern the most likely BPS field configurations inherent to the proposed theory. In Sec. IV, one analyzes both the excitation spectrum and the scattering phenomena of the solutions. Finally, we present our remarks and conclusions in Sec. V.
@@ II.1 Modifying the $\phi^{4}$ model
To continue our study, we need to choose both the superpotential $W(\phi)$ and the generalizing function $f(\phi)$ to define a BPS potential (6) that preserves the $\mathds{Z}_{2}$ [^3] A system possesses a $\mathds{Z}_{2}$ symmetry when the action is invariant under the transformation $\phi\rightarrow-\phi$. symmetry [6, 36] but at the same time able to promote the spontaneous symmetry breaking of it. This way, for our purpose, we use a $\phi^{4}$ model [37, 38] defined by the following superpotential,
@@
III The Differential Configurational Complexity (DCC)
[…]
Now, let us consider the numerical solutions depicted in Figs. 2 and 3, and employing a numerical approach [^5] The numerical method applied was the interpolation method. This method is used with steps of $10^{-4}$ for the range $[-1000,1000]$., we compute the integrals (20) and (22). We expose the numerical solutions obtained in Figs. 5(a) and 5(b).
(a)                                                                 (b)
IV Excitation spectra and scattering of structures
@@ IV.1 Excitation spectra
We display the stability potential (30) in Figs. 7, 9, and 11.
Utilizing the finite element method, we investigate the model eigenstates [19, 43]. We first illustrate in Figs. 7 and 8 the stability potential, and corresponding eigenfunctions of the case $m=0$ (i.e., those of the standard $\phi^{4}$-theory). In this scenario, the model admits a translational mode (or zero mode, i.e., the eigenfunction with null eigenvalue, $\omega=0$) and one vibrational eigenstate ($\omega^{2}>0$). Besides, one notes through numerical inspection that this spectrum remains unchanged for sufficiently small $m$ values, but starting from sufficiently larger $m$ values, one translational mode and several vibrational modes emerge. The stability potential, translational mode, and first vibrational modes, along with the associated eigenvalues, are depicted in Figs. 9[(a) and (b)], 10, and 11[(a) and (b)], respectively. Notably, the presence of the translational and vibrational modes suggests the occurrence of resonances in the scattering of configurations with opposing topological charges.
@@
(a)                                                             (b)
Figure 9: (a) Stability potential $U(x)$ (30) keeping $\lambda=1$ and $m$ varying. (b) Stability potential $U(x)$ (30) keeping $m=0.45$ and $\lambda$ varying.
Figure 10: Translational mode, i.e., eigenfunctions $\eta(x)$ vs. $x$ for the eigenvalue null.
@@
Figure 11: (a) Eigenfunctions of Eq. (25) representing the ground states (first vibrational modes) for some values of $m$ and $\lambda=1$. (b) Eigenfunctions of Eq. (25) representing the ground states (first vibrational modes) for some values of $\lambda$ and $m=0.45$
@@
The results show the existence of translational and vibrational modes, indicating that when configurations tend towards a compact-like profile, the resonance phenomenon will occur in kink-antikink scattering. Similarly, one emphasizes that experimental studies on kink collisions have also observed resonance phenomena even in systems without vibrational modes [19]. Furthermore, we are aware of two possible collision types for the system: the first produces energy emission with the presence of the resonance phenomenon, and the second is a collision in which the kink interactions form bions along the time evolution and energy dissipation [8, 55].
@@ IV.2 Scattering of structures
Figures 14 and 15 show the analysis of the collisions of the kink-antikink structures for several initial velocity values. One notes that also, throughout the scattering, kink/antikink solutions behave like compacted configurations when $m\to 0.5$ and $\lambda\to\infty$. Besides, it exhibits critical velocities ($v_{\text{cr}}$) that alter the scattering profile. For instance, for $v_{\text{in}}<v_{\text{cr}}^{(1)}\simeq 0.15$, it stays evident that during the temporal evolution, the kink-antikink collide, annihilating each other and radiating energy. Conversely, one obtains an inelastic collision for an initial velocity whose value belongs to the interval <0.15,0.25]<0.15,0.25]< 0.15 , 0.25 ]. On the other hand, for $v_{\text{in}}>0.25$, the compacton-like configurations, here generated, experiment quasi-elastic collisions; see Fig. 16. Thus, the compacton-like configurations in the external scattering process support both inelastic and quasi-elastic collisions, and they manifest as solitary waves with physical characteristics akin to kinks.
@@
By looking closest, the numerical results shown by Figs. 14, 15, and 16, it is noted that for initial velocities smaller than $0.25$, the $K\bar{K}$ pair interacts forming oscillations (bions [^6] Bions arise during the collision of kinks and have interesting characteristics, such as topological charges, which are a combination of the topological charges of the original kinks. One can interpret the bions as a pair of “temporary kinks” that form during a collision.).For instance, when $v_{\text{in}}=0.10$, the first oscillation emerges. These oscillations increase as the velocity increases, regardless of the values of $m$ or $\lambda$. However, upon reaching the value of $v_{\text{in}}=0.25$, as time evolution, only three oscillations will occur, and then the $K\bar{K}$ pair scatters with the resonance phenomenon (oscillation amplitudes between $0.8$ and $1.2$) present for a time range greater than $200$. Finally, one notes that when $v_{\text{in}}\geq 0.35$, regardless of the value of $\lambda$, there will be a quasi-elastic scattering process with extreme regions of the $K\bar{K}$ pair undergoing oscillations (resonance) between 00 and $1.2$
@@ V Final remarks
We continue our study by applying the DCC technique to determine the most suitable values for the $m$ parameter of the function $f(\phi)=\cos(m\pi\phi)$ and for the parameter $\lambda$. Thus, through numerical inspection, the DCC leads us to conclude that field configurations with lower complexity are more likely to emerge when $\lambda\simeq 300$ and $m\to 0.5$, i.e., compacton-like structures. It is a significant result because, despite infinite degeneracy, values of $m$ and $\lambda$ exist, describing the best configuration possible for the system.
We conclude our study by analyzing the excitation spectrum of the model for fixed $\nu$ and $\lambda$. Thus, we find a discrete spectrum with translational and vibrational modes. We note that the resonance phenomena are guaranteed to occur during the scattering process of configurations with opposite topological charges. This way, the study reveals that the scattering of the compacton-like solutions presents the following features depending on the value of the initial velocity ($v_{\text{in}}$): for $v_{\text{in}}\simeq 0.15$ they collide and annihilate each other by radiating energy. On the other hand, for an initial velocity occurring in the interval $[0.15,0.25]$, an inelastic collision happens. These results reveal that the compacton-like configurations behave as solitonic waves with physical aspects similar to the ones arising in the kink/antikink collision processes. Furthermore, the compacton-like solutions collide quasi-elastically when the initial velocity is $v_{\text{in}}>0.25$.